PartA Determine the inductance L of a 0.40-m-long air-filled solenoid 2.7 cm in diameter containing 8100 loops. Express your answer
Answer #1

Inductance of a solenoid is given by:

L = u0*N^2*A/L

N = 8100 loops

L = 0.40 m

A = pi*d^2/4 = pi*0.027^2/4

A = 5.72*10^-4 m^2

Using given values:

L = 4*pi*10^-7*8100^2*5.72*10^-4/0.4

L = 0.1179 = 0.12 H